OpenStudy (anonymous):

Theo wants to use a cookie recipe that makes 36 cookies, but he wants to reduce the number if cookies to 24. If the recipe specifies 2 cups of sugar, how much sugar should he use?

OpenStudy (anonymous):

btw this is dealing with proportions :I

OpenStudy (anonymous):

He shouldn't use any sugar. That stuff is dangerous. It will harm you. But if you wanna get diabetes that badly then \[ \frac{2}{36}=\frac{x}{24} \]

OpenStudy (anonymous):

what does x equals than? I got 1.5, but I think i got it wrong

OpenStudy (anonymous):

If they are proportional, then the ratios have to be equal, hense the equation.

OpenStudy (anonymous):

I got 1.5 of a cup

OpenStudy (anonymous):

You got it wrong.

OpenStudy (anonymous):

let me redo the math.

OpenStudy (anonymous):

\[ \frac{2}{36}=\frac{x}{24}\implies 36x = 2\cdot 24\implies x=\frac{48}{36}=\frac 43 \]

OpenStudy (anonymous):

I just need to know how much sugar he needs to use :I

OpenStudy (anonymous):

I said \(x=4/3\)